Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.)
3 tan^3theta = tantheta

Answer by lwsshak3(11628)

Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.)
3 tan^3theta = tantheta
**
use x for theta
3tan^3x=tanx
3tan^3x-tanx=0
tanx(3tan^2x-1)=0
tanx=0
x=0+2πk, π+2πk, k=integer
or
3tan^2x-1=0
tan^2x=1/3
tanx=1/√3
x=π/6+2πk, 7π/6+2πk, k=integer